Find the pH of a 0.0150M H2SO4 solution (Ka=0.012)
Answer #1

               H2SO4 is strong acid

   H2So4 ----------> H^+ (aq) + HSO4^-

I        0.015                  0                   0

C      -x                      +x                    +x

E     0.015-x              +x                     +x

      ka     =   [H^+][HSO4^-]/[H2SO4]

    0.012   = x*x/0.015-x

   0.012*(0.015-x) = x^2

    x = 0.0087

   [H^+]   = x = 0.0087 M

PH   = -log[H^+]

          = -log0.0087

          = 2.0604 >>>>>answer